You own a call option on the stock with current price So-4, the up factor, u = 2, the down factor d-1/2, the risk-free interest rater-1/4 and the strike price K 5. You paid the risk-neutral price of $1.20 for this option, and you want to hedge your position (i.e. reduce your risk) so that you end up with $1.50, (as if you had invested $1.2 at the risk free rate) regardless of whether H or T occurs. Assume that you cannot sell the option, and you have no cash. You may only buy (or short sell) shares of stock or bond (borrow or invest at the risk-free interest rate), construct a portfolio to ensure you have $1.50 att
Answer #1

Let's create a replicating portfolio

  • Using the call option that you have,
  • Short N numbers of underlying stock, and
  • Lend B at risk free interest rate

So, the value of this portfolio when the stock is in up state can be calculated as shown below:

Price of the stock in the up state = Su = S0 x u = 4 x 2 = 8

  • Value of the call option = max (Su - K, 0) = max (8 - 5, 0) = 3
  • Maturity value of the amount lent = B x (1 + r) = B x (1 + 1/4) = 1.25B.
  • Value of N numbers of underlying stock short sold = -N x Su = -N x 8 = -8N (The negative sign signifies a payout from us on account of closing the short sale position)

Hence value of the portfolio in up state = 3 + 1.25B - 8N

This should be equal to $ 1.50. Hence, we get first equation as : 3 + 1.25B - 8N = 1.50

Or, 8N - 1.25B = 1.50

So, the value of this portfolio when the stock is in down state can be calculated as shown below:

Price of the stock in the down state = Sd = S0 x d = 4 x 1/2 = 2

  • Value of the call option = max (Sd - K, 0) = max (2 - 5, 0) = 0
  • Maturity value of the amount lent = B x (1 + r) = B x (1 + 1/4) = 1.25B.
  • Value of N numbers of underlying stock short sold = -N x Sd = -N x 2 = -2N

Hence value of the portfolio in down state = 0 + 1.25B - 2N = 1.25B - 2N

This should be equal to $ 1.50. Hence, we get second equation as : 1.25B - 2N = 1.50

Add the two equations

(8N - 1.25B) + (1.25B - 2N) = 1.50 + 1.50

Hence, 6N = 3. Hence N = 3/6 = 0.5

Further, 8N - 1.25B = 1.50

Hence, 8 x 0.5 - 1.25B = 4 - 1.25B = 1.50

Hence, B = (4 - 1.50) / 1.25 = 2

Hence the strategy should be:

  1. Short sell half number of stocks and
  2. Lend $ 2 at risk free rate

Initial investment required = proceeds from short sale of 0.5 stock - amount lent = 0.5 x S0 - 2 = 0.5 x 4 - 2 = 0. No cash (No initial investment) is required for this strategy.